OpenStudy (anonymous):

find the new function of f(x)=ln x if f(x) is reflected over the line y=x, then over y axis, and finally over the x-axis.

OpenStudy (anonymous):

lol, it is f(x)=e^x

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

how did you get thta?

OpenStudy (anonymous):

when you reflect f(x)=ln(x) over line y=x (which is a diagonal line running from Q1 to Q3, the reflected graph is the graph of e^x, actually e^x is the inverse of ln(x) and the prove is that ln(e) = 1

OpenStudy (anonymous):

if you have calculator, check ln(e)

OpenStudy (anonymous):

wait wait, i didnt read the question completely

OpenStudy (anonymous):

same, as i said b4, the question continuous on saying, "if f(x) is reflected over the line y=x, then over y axis, and finally over the x-axis." if you do all those reflections in your mind, you end up with f(x)=ln(x) = and the answer is ln(x)

OpenStudy (anonymous):

yes correct it is ln(x) back again, because you diagonally reflect ln(x), reflect around y-axis, then again around x-axis, you end up with same function's graph, so the new function is the old function is ln(x), sorry for not reading the question completely :)

OpenStudy (anonymous):

haha it's cool. thank you for your help! :)